# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

### v1

@@ 3.3 Main Results
See Appendix A.3. ∎
Remark 3.8.
Theorem 3.7 can be concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ into Zhao et al. (2024, Theorem 4.3). This implies that the sample complexity lower bound in Theorem 3.7 also holds for KL-regularized dueling contextual bandit with all-policy concentrability.
Remark 3.9.
Theorem 3.7 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandit requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output a $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 3.5, indicating that KL-PCDB is nearly optimal.
3.4 Proof of Theorem 3.5
$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (3.3)
Here, $\Gamma_{n}$ is defined in (3.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (3.4)
where $\Phi_{b}(s,a)=\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 2, after adjusted by some bias function $b$, is close to the true function $f^{*}$. The following lemma shows that this event holds with high probability. Please refer to Appendix B.1 for the proof.
Lemma 3.10.
For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
@@
$f_{\gamma}(s,a)\coloneqq\gamma({\widehat{f}}(s,a)-b(s))+(1-\gamma)f^{*}(s,a).$
@@
$\pi_{\gamma}(\cdot|\cdot)=\pi_{f_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta f_{\gamma}(\cdot,\cdot)\big{)},$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]},$
where $b(\cdot)$ is defined in (3.4). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma 3.11.
@@
For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)$ in this proof. Then on event $\mathcal{E}(\delta)$, we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$. Taking derivatives of $G$ w.r.t., $\gamma$ directly, we conclude that for all $\gamma\in[0,1]$,
@@
Proposition 3.12.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $f\in\mathcal{F}$ we have $J(\pi_{f})=J(\pi_{f-b})$, where $(f-b)(s,a)=f(s,a)-b(s)$.
Please see Appendix B.2 for the proof of Proposition 3.12. Now we are ready to prove Theorem 3.5.
Proof of Theorem 3.5.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition 3.12, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}})$
$=J(\pi^{*})-J(\pi_{{\widehat{f}}-b}).$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}-b})$
$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]}$
$=\eta G(\gamma),$ (3.5)
@@
where the inequality holds due to Lemma 2.17. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{f}}(s,a)-b(s)\leq{f^{*}}(s,a)$. Together with Lemma 3.11, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$ (3.6)
$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\Big{]}$
$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\Gamma_{n}^{2}(s,a)$
$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (3.7)
@@
where the inequality holds due to the definition of $\mathcal{E}(\delta)$. Plugging (3.7) into (3.5), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$. By Lemma 3.10, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof. ∎
4 Conclusions and Future Work
In this work, we settled the sample complexity of KL-regularized offline contextual bandits. We first propose KL-PCB, which follows the principle of pessimism for reward estimation and yields the optimal policy accordingly. Our analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified covariance-based observation, which in turn enables a neat refinement of a mean-value-type argument the its extreme point, showing that KL-PCB achieves a nearly optimal sample complexity under single-policy concentrability. The pair of novel techniques are decoupled from tricky algorithmic tweaks, and thus might be of independent interest. We then extended KL-PCB to offline KL-regularized contextual dueling bandit problem and introduced KL-PCDB, the analysis of which manifests the near-optimality of KL-PCDB under single-policy concentration. Currently, our focus is solely on the reverse-KL-regularized objective. It would be an interesting future direction to study more general regularization techniques like $f$-divergence (Csiszár, 1967) regularization, which includes both forward and reverse KL regularizations.
Appendix A Missing Proofs from Section 2
A.1 Proof of Lemma 2.13
We first provide the following lemmas.
Lemma A.1 (Zhao et al. 2024, Lemma C.1).
For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d. from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $f_{1}$ and $f_{2}$ we have
$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}f_{1}(s_{i},a_{i})-f_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
where $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of $\mathcal{F}$.
Lemma A.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{f}}$ be the least square estimator of $f^{*}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
$\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
Proof of Lemma 2.13.
We have the following inequality
$\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\leq\sup_{f_{1},f_{2}\in\mathcal{F}}\frac{\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$=D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]},$ (A.1)
where the inequality holds by taking supremum to $f_{1},f_{2}\in\mathcal{F}$. Now we have
$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$=\frac{128}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (A.2)
where the first inequality holds due to Lemma A.1 and second holds due to Lemma A.2. Plugging (A.1) into (A.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
A.2 Proof of Lemma 2.17
$J(\pi^{*})-J(\pi_{f})$
$\quad=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}f^{*}(s,a)-\eta^{-1}\log\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}-\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}f^{*}(s,a)-\frac{1}{\eta}\log\frac{\pi_{f}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi^{*}(a|s)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi_{f}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f^{*}}(s)\big{]}-\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f}(s)\big{]}-\mathbb{E}_{s\sim\rho}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\big{(}f^{*}(s,a)-f(s,a)\big{)}\bigg{]},$
where for all $f\in\mathcal{F}$ we define $Z_{f}(\cdot)$ as follows,
$Z_{f}(\cdot)\coloneqq\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a|\cdot)\exp\big{(}\eta f(\cdot,a)\big{)}.$
We further denote $\Delta(s,a)=f(s,a)-f^{*}(s,a)$ and $H_{s}(f)=\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)$. It worth noticing that $\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]=J(\pi^{*})-J(\pi_{f})$. Now we take derivative of $H$ with respect to $\Delta(s,a)$,
$\frac{\partial H_{s}(f)}{\partial\Delta(s,a)}$ $=\frac{\partial}{\partial\Delta(s,a)}\bigg{[}\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)\bigg{]}$
$=\frac{1}{Z_{f}(s)}\cdot\pi^{\mathsf{ref}}(a|s)\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\cdot\eta-\eta\cdot\pi_{f}(a|s)$
$\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl{[}\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\bigr{]}^{2}}{[Z_{f}(s)]^{2}}$
$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot f(s,a^{\prime})\bigr{)}}{Z_{f}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}$
$=-\eta^{2}\pi_{f}(a|s)\Delta(s,a)+\eta^{2}[\pi_{f}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{f}(a^{\prime}|x)\pi_{f}(a|s)\Delta(s,a^{\prime}).$
Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ such that
$H_{s}(f)-H_{s}(f^{*})$ $=-\eta^{2}\gamma\sum_{a\in\mathcal{A}}\pi_{f_{\gamma}}(a|s)\Delta(s,a)^{2}+\gamma\eta^{2}\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{f_{\gamma}}(a_{1}|x)\pi_{f_{\gamma}}(a_{2}|x)\Delta(s,a_{1})\Delta(s,a_{2})$
$=-\eta^{2}\gamma\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]}+\gamma\eta^{2}\Big{(}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}\big{]}\Big{)}^{2}$
$\geq-\eta^{2}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
$J(\pi^{*})-J(\pi_{f})$ $=\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]$
$\leq\eta\mathbb{E}_{(s,a)\sim\rho\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
which concludes the proof.
A.3 Proof of Theorem 2.10
Proof of Theorem 2.10.
We consider the bandit instance $\mathcal{K}$ and corresponding function class $\mathcal{F}$ constructed in the proof of Zhao et al. (2024, Theorem 3.1). Since the constructed $\pi^{\mathsf{ref}}(\cdot|s)$ in $\mathcal{K}$ is a uniform distribution for any $s\in{\mathcal{S}}$, the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ defined in Assumption 2.6 satisfies $D^{2}=O(1)$, which indicates that the single-policy concentrability coefficient $D_{\pi^{*}}^{2}$ defined in Assumption 2.7 satisfies $D_{\pi^{*}}^{2}\leq D^{2}=O(1)$. Thus, the proof of Zhao et al. (2024, Theorem 3.1) implies that any Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy. ∎
Appendix B Missing Proofs from Section 3
B.1 Proof of Lemma 3.10
The following on-poicy concentration bound is the prerequisite.
Lemma B.1 (Zhao et al. 2024, Lemma D.4).
Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 3.1. Suppose ${\bar{f}}$ is the result of MLE estimation of Algorithm 2, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
$\mathbb{E}_{s,a\sim\rho\times\pi}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)-b(s)\big{)}^{2}\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$
Proof of Lemma 3.10.
From Lemma B.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (B.1)
It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
$D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}\geq\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}.$ (B.2)
Substituting (B.1) into (B.2), we have
$\inf_{b}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}$
$\quad=\inf_{b}\frac{\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$
where the first inequality holds due to the definition of $D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma B.1. ∎
B.2 Proof of Proposition 3.12
Proof of Proposition 3.12.
$\pi_{f}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta f(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta f(s,a^{\prime})\big{)}}$

### v2

@@ Appendix E Missing Proof from Appendix D
From Lemma F.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (E.3)
It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}\geq\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}.$ (E.4)
Substituting (E.3) into (E.4), we have
$\inf_{b}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}$ (E.5)
$\quad=\inf_{b}\frac{\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}$
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ (E.6)
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$ (E.7)
where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma F.1. ∎
@@
$g_{\gamma}(s,a)\coloneqq\gamma({\widehat{g}}(s,a)-b(s))+(1-\gamma){g^{*}}(s,a).$
@@
$\pi_{\gamma}(\cdot|\cdot)=\pi_{g_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta g_{\gamma}(\cdot,\cdot)\big{)},$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\big{]},$
where $b(\cdot)$ is defined in (E.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma E.2.
@@
For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)$ in this proof. Then on event $\mathcal{E}(\delta)$, we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$. Taking derivatives of $G$ w.r.t., $\gamma$ directly, we conclude that for all $\gamma\in[0,1]$,
@@
Proposition E.3.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where $(g-b)(s,a)=g(s,a)-b(s)$.
Proof.
$\pi_{g}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta g(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta g(s,a^{\prime})\big{)}}$
[…]
$=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta[g(s,a)-b(s)]\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta[g(s,a^{\prime})-b(s)]\big{)}}$
$=\pi_{g-b}(a|s),$
which indicates that $\pi_{g}=\pi_{g-b}$. This immediately leads to $J(\pi_{g})=J(\pi_{g-b})$. ∎
Now we are ready to prove Theorem D.1.
Proof of Theorem D.1.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition E.3, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}})$
$=J(\pi^{*})-J(\pi_{{\widehat{g}}-b}).$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}-b})$
$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\big{]}$
$=\eta G(\gamma),$ (E.8)
@@
where the inequality holds due to Lemma 2.16. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{g}}(s,a)-b(s)\leq{{g^{*}}}(s,a)$. Together with Lemma E.2, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$
$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\Big{]}$
$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\big{[}\Gamma_{n}^{2}(s,a)\big{]}$
$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big{)},$ (E.9)
@@
where the inequality holds due to the definition of $\mathcal{E}(\delta)$. Plugging (E.9) into (E.8), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$. By Lemma E.1, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof. ∎
E.2 Proof of Theorem D.3
Proof of Theorem D.3.
The proof is similar to the proof of Theorem 2.10. Consider the following family of contextual dueling bandit instances with $S\coloneqq|{\mathcal{S}}|,A\coloneqq|\mathcal{A}|<\infty$ and reward in some function class $\mathcal{G}$.
$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (E.10)
Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq 2C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (E.11)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, the above bound yields the desired result.